The length of a rocket is measured to be 6.0 m by an observer on Earth as it moves
with respect to the Earth at a speed of 0.8c. What is the length of the rocket as
measured by an astronaut inside the rocket?
1
Expert's answer
2019-01-09T10:08:11-0500

l_0=l/√(1-v^2/c^2 )=6/√(1-(0.64c^2)/c^2 )=6/0.6=10 meter